OpenStudy (babyslapmafro):

Please help me get started on the following integration (click to see)...

OpenStudy (babyslapmafro):

\[\int\limits_{}^{}\frac{ 1 }{ \cos ^{2}x }\]

OpenStudy (anonymous):

Well we do know that \(\frac{d}{dx}\tan(x)=\frac{1}{\cos^2(x)}\) So we can use the relationship of: \[F(x)=\int{f(x)}\] \[\frac{d}{dx}F(x)=f(x)\] \(F(x)=\tan(x)\) \(f(x)=\frac{1}{\cos^2(x)}\) \[\frac{d}{dx}\tan(x)=\frac{1}{\cos^2(x)}\] \[\therefore:\phantom{spc}\tan(x)=\int{\frac{1}{\cos^2(x)}}\]